OpenStudy (anonymous):

7k+12=8-9k show work please. :)

OpenStudy (kropot72):

Add 9k to both sides: 7k + 9k + 12 = 8 - 9k +9k 16k +12 = 8 Add -12 to both sides: 16k + 12 - 12 = 8 - 12 16k = -4 k = -4/16 = -(1/4) or -0.25

OpenStudy (anonymous):

7k+12 = 8-9k 7k + 9k + 12 = 8 16k + 12 = 8 16k = 8-12 16k = -4 k = -4/16 k = -1/4